OpenStudy (anonymous):

Find the x and y intercepts of the graph line 1) y=3x+2 2) 3x+4y=12

OpenStudy (marcelie):

x intercept plugin 0 for y y inetrcept plugin 0 for x and make the 2nd equation into y=mx+b

Directrix (directrix):

1) y=3x+2 --------- The x-intercept is on the x-axis and has a y-value of 0. y=3x+2 0 = 3x + 2 3x = -2 x = -2/3 x-intercept (-2/3 , 0) ==== The y-intercept is on the y-axis and has x-coordinate 0. y=3x+2 y = 3*0 +2 y = 2 Coordinates of y-intercept (0,2) ========== Use those as examples to find the x-and y-intercepts of 2) 3x+4y=12 @Alainambri